OpenStudy (anonymous):

write an equation of a sine function with amplitude 3, period 3pi/2 and phase shift pi/4

OpenStudy (anonymous):

is it y=3sin(3x/2-pi/4)?

OpenStudy (anonymous):

3sin((4/3)t+(pi/4))

OpenStudy (anonymous):

what about -3sin(4x/3-pi/4)?

OpenStudy (anonymous):

general equation of a simple harmonic oscillator is Asin(wt). where A is the amplitude and w is the angular frequency.. amd 2pi/w is the time period...

OpenStudy (anonymous):

the equation you are giving should be a function of time and not distance... and by adding a minus sign you are just changing its phase by pi.

OpenStudy (anonymous):

so my answer is wrong?

OpenStudy (anonymous):

a. y = -3 sin (3x/2 - 3π/8) b. y = 3 sin (4x/3 - π/3) c. y = -3 sin (4x/3 - π/4) d. y = 3 sin (3x/2 - π/4)

OpenStudy (anonymous):

those are my choices @AERONIK

OpenStudy (anonymous):

you got me wrong friend, your choices are definitely correct....

OpenStudy (anonymous):

wait i'm confused

zepdrix (zepdrix):

\[\huge y=a \sin(bx-c)+d\]\[a=amplitude\]\[\frac{2\pi}{b}=period\]\[c=phase \; shift\]

zepdrix (zepdrix):

Let's figure out our b term. \[\large \frac{2\pi}{b}=\frac{3\pi}{2}\]\[\large b=\frac{4}{3}\] I hope I calculated that correctly hehe

zepdrix (zepdrix):

\[\large y=3\sin\left(\frac{4}{3}x-\frac{\pi}{4}\right)\]

zepdrix (zepdrix):

Mmmmm I'm not sure if I did that correctly, was I suppose to factor the 4/3 into the pi/4? I forget...

OpenStudy (anonymous):

thats not an option :(

zepdrix (zepdrix):

Hmm

OpenStudy (anonymous):

wait yes i think so

zepdrix (zepdrix):

Yah I think the b is suppose to be like this... \[\huge y=a \sin (b(x-c))+d\]

zepdrix (zepdrix):

Which IS one of your options, if you distribute the 4/3 to the pi/4 term.

zepdrix (zepdrix):

Hmm how did you arrive at your answer? :o

OpenStudy (anonymous):

i think its b. y = 3 sin (4x/3 - π/3)

zepdrix (zepdrix):

Hmmm I think so too :O